OpenStudy (anonymous):

how do i solve 5x-7y=9 4x+3y=33

OpenStudy (anonymous):

this with it multiple choice

OpenStudy (anonymous):

I'm assuming that those are two separate equations that should look like this: \(5x - 7y = 9\) and \(4x +3y = 33\) In that case, choose one equation (it doesn't matter which one) and solve for a variable (again, it doesn't matter which one). Then, substitute the resulting expression into the other equation in place of the variable that you have chosen. I'll provide an example: \(2x – 3y = –2\) and \(4x + y = 24\) I'm going to choose the second equation and solve for y. \(4x + y = 24 \) First, isolate y by subtracting both sides by 4x: \(y = 24-4x\) If y had a coefficient(other than 1) you would divide both sides by that number. It doesn't so we are done with this step. Now, Substitute \(24-4x\) for y in the other equation: \(2x-3(24-4x) = -2\) And solve this equation using Order of Operations. Parenthesis: \(2x-72+12x = -2\) Combine like terms by adding 72 to each side then combine the x terms: \(14x = 70\) Divide both sides by 14 to isolate the x: \(x = 5\) Now, substitute 5 into each equation and solve for y: (hint: you should get the same answer on each equation - this is how you know you are correct.) \(2x – 3y = –2\) and \(4x + y = 24\) \(2(5) – 3y = –2\) and \(4(5) + y = 24\) \(10 – 3y = –2\) and \(20 + y = 24\) \(-3y = -12\) and \(y = 4\) \(y = 4\) If there is anything that is unclear, please ask.

OpenStudy (texaschic101):

good job gypsy :)

OpenStudy (anonymous):

Thanks Texaschic.

OpenStudy (anonymous):

i dont see how you got 2x at the beginning of the problem @gypsy1274

OpenStudy (anonymous):

@papsmurf I don't like providing answers, so I created a different example to explain the steps. The original poster was already offline when I answered this question.

OpenStudy (anonymous):

ohh ok